# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

### v2

@@ 4.2. Comparison to the previous state of the art
By contrast, the functions in the CM suite have numerous dependencies, often use external
code and are generally unannotated.
Across the board, CoverUp yields higher coverage than CodaMosa,
whether measured over the entire suite or on a module-by-module basis.
Figure 11. [RQ1] CoverUp yields higher overall and median per-module coverage:CoverUp yields higher or equal coverage than MuTAP,
whether measured over the entire suite or on a module-by-module basis.
The graph shows the difference between the lines + branches coverage achieved by
CoverUp and CodaMosa for the various benchmark modules in the CM benchmark suite,
sorted by percentage of coverage increase.
Green bars indicate modules where CoverUp achieved a higher coverage.
Figure 13. [RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage
on all metrics, showing that its effectiveness is not limited to
the modules selected for the CM suite.
Figure 14. [RQ3] Contribution of iterative coverage-guided approach:
nearly 40% of the successes were obtained
by iterating and providing coverage and error feedback,
demonstrating its importance.
We next evaluate the performance of CoverUp on code that the
original Pynguin already performed well.
Figure 14 shows the results.
Given CoverUp’s 100% median per-module coverage and near 100% overall coverage
results, we conclude that CoverUp also performs well on such code.
[RQ1] Summary:CoverUp achieves significantly higher coverage than
both CodaMosa and MuTAP, outperforming the state of the art.
4.3. How effective is CoverUp compared to simply prompting an LLM for tests?
Given many LLMs’ near-human performance on various tasks, including software
engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how
much CoverUp contributes to performance.
To address this question, we compare CoverUp to an ablated LLM-only
version that utilizes a nearly identical prompt, but which lacks all other
CoverUp functionality (coverage information, precomputed computed
import statements, tool function for additional context
and continued chats).
As Figure 15 shows,
CoverUp achieves substantially higher line, branch and combined line and branch coverage
than CoverUp (ablated) both measuring over the entire benchmark code base
and on a per-module basis.
Across the entire benchmark suite, it achieves 64% (vs. 39%) line coverage,
49% (vs. 19%) branch coverage, and 60% (vs. 34%) line+branch coverage;
on a per-module basis, it achieves 82% (vs. 44%) line coverage,
67% (vs. 7%) branch coverage, and 80% (vs. 39%) line+branch coverage.
@@ 4.3. How effective is CoverUp compared to simply prompting an LLM for tests?
Across the board, CoverUp yields higher coverage in comparison to
the ablated (LLM only) version, whether measured over the entire suite
or on a module-by-module basis.
[RQ2] Summary:CoverUp contributes significantly to overall performance;
its performance is not just due to the LLM employed.
4.4. How effective are CoverUp’s continued dialogues?
To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite,
identifying each successful test (i.e., which passes and improves coverage) that
was generated immediately upon the first prompt or after continuing the chat with a second
or third prompt
(by default, and also in our evaluation, CoverUp continues the conversation for at most
two additional prompts).
Figure 14 shows the percentage of successes after
the first, second and third prompts.
As the figure shows, while the success rate decreases with each iteration, about
40% of the successes were obtained through iterative refinement of the prompt,
demonstrating its importance.
[RQ3] Summary: continuing the chat contributes almost half its successes,
demonstrating its effectiveness.
5. Threats to Validity
Benchmark selection
We utilize the CM and PY benchmark suites to evaluate CoverUp’s
performance on both challenging and less challenging code, as well
as to facilitate including Codex-based results;
we utilize the MT benchmark suite because it is supported by MuTAP
and it, too, facilitates including Codex-based results.
While so far our experience executing CoverUp to generate tests
for other software has yielded similarly high coverage, selecting
a different set of benchmarks could yield different results.
Execution environment
CodaMosa’s original evaluation environment failed to install a number of
Python modules that are prerequisites for the applications used as benchmarks.
In an effort to replicate the original conditions as closely as possible,
we ignored these failures as well.
It seems likely that both CodaMosa and CoverUp would be better able
to generate tests if these modules were not missing.
LLM model dependency
CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo
and GPT-4o models.
While CoverUp’s test execution and coverage measurement portions
are independent of the LLM
and, as Section 4.3 shows, CoverUp significantly
outperforms an ablated version of itself that just relies on the LLM’s capabilities,
its ultimate performance naturally depends on the model’s ability to
interpret its prompts and generate tests as requested.
6. Conclusion
This paper presents CoverUp, a novel approach for guiding
LLM-based test generation with coverage analysis.
We show that coupling coverage information with prompting, and iteratively
refining prompts with updated coverage information to focus the LLM’s attention
on improving uncovered code, is effective at generating test suites with
higher coverage than past approaches.
Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and
far outperforms itself when ablated to simply rely on LLM performance.
@@ 6. Conclusion
metric
@@
l+b
module median
overall
Table 3. [RQ1] CoverUp’s performance is not limited to “challenging” code: results for PY suite
prompt
contribution
initial
second
third
Table 4. [RQ3] Contribution of iterative coverage-guided approach to successes.

### v3

@@ 4.1. Experimental Setup
By contrast, the functions in the CM suite have numerous dependencies, often use external code, and are generally unannotated.
[RQ1] Summary:CoverUp achieves significantly higher coverage than both CodaMosa and MuTAP, outperforming the state of the art.
Overall Coverage
Median Per-Module Coverage
Test Generator
Line
Branch
Line + Branch
Line
Branch
Line + Branch
CoverUp
CodaMosa (codex)
CodaMosa (gpt4o)
CoverUp (ablated)
CoverUpon PY suite
Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and far outperforms itself when ablated to simply rely on LLM performance (top).
Bottom line: as the results for the PY suite show, CoverUp’s performance is not limited to “challenging” code.
@@
Across the board, CoverUp yields higher coverage in comparison to the ablated (LLM only) version, whether measured over the entire suite or on a module-by-module basis.
4.3. [RQ2] How effective is CoverUp compared to simply prompting an LLM for tests?
Given many LLMs’ near-human performance on various tasks, including software engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how much CoverUp contributes to performance.
To address this question, we compare CoverUp to an ablated LLM-only version that utilizes a nearly identical prompt but which lacks all other CoverUp functionality (coverage information, computed import statements, get_info tool function and continued chats).
Rather than ask for tests for the uncovered code, it asks instead for tests “that execute all lines and branches”.
As Figure 16 shows, CoverUp achieves substantially higher line, branch, and combined line and branch coverage than CoverUp (ablated), both measured over the entire benchmark code base and on a per-module basis.
Across the entire benchmark suite, it achieves 64% (vs. 39%) line coverage, 49% (vs. 19%) branch coverage, and 60% (vs. 34%) line+branch coverage; on a per-module basis, it achieves 82% (vs. 44%) line coverage, 67% (vs. 7%) branch coverage, and 80% (vs. 39%) line+branch coverage.
[RQ2] Summary:CoverUp contributes significantly to overall performance; its performance is not just due to the LLM employed.
@@
4.4. [RQ3] How effective are CoverUp’s continued dialogues?
To investigate RQ3, we process CoverUp’s logs generated while evaluated on the CM suite, identifying each successful test (i.e., which passes and improves coverage) that was generated immediately upon the first prompt or after continuing the chat with a second or third prompt (by default, and also in our evaluation, CoverUp continues the conversation for at most two additional prompts).
We observe that 60.3% of successes result from the first prompt, 27.2% from the second, and 12.4% from the third.
While the success rate decreases with each iteration, approximately 40% of successes were achieved through iterative refinement of the prompt, highlighting its importance.
[RQ3] Summary: Continuing the chat contributes about 40% of its successes, demonstrating its effectiveness.
4.5. [RQ4] How does the cost of running CoverUp compare to CodaMosa?
Examining the logs from runs on the CM suite, we assemble Table 3, which shows the number of prompts, completions, tokens, running time, and approximate cost incurred.
Test Generator
Prompts
P. Tokens
Completions
C. Tokens
Time (h)
Cost (US$)
CoverUp
28,690
58,081,908
28,654
7,257,400
4.0
399
CoverUp (ablated)
9,224
3,501,827
9,188
3,859,444
1.7
75
CodaMosa (gpt4o)
30,368
39,508,559
29,136
4,504,329
71.0
265
Table 3. [RQ4] Cost of running CoverUp and CodaMosa on the CM suite:CoverUp runs roughly 18 faster while achieving higher coverage
than CodaMosa, but uses 48% more tokens.
We observe that CoverUp runs roughly 18 times faster (4 vs. 71 hours) while achieving higher coverage than CodaMosa, but using 48% more tokens.
Although this time comparison can provide insight into each method’s applicability for a given context, it is important to understand its limitations.
CoverUp and CodaMosa utilize both local and cloud-based computing resources; we describe the local system in Section 4.1, but OpenAI does not disclose details of their computing environment, whose availability may vary over time.
CoverUp’s implementation is both parallelized and asynchronous.
Its running time is primarily gated by the rate limits imposed by OpenAI (see Section 3.7), their response times, and the time required to execute the generated tests.
In contrast, CodaMosa is a sequential process that runs for 10 minutes per module.
While this duration is configurable, we follow the setting used in the CodaMosa paper (Lemieux et al., 2023).
[RQ4] Summary: As deployed, CoverUp runs 18 times faster while achieving higher coverage than CodaMosa, but using almost 50% more tokens.
4.6. [RQ5] How important are CoverUp’s components to its performance?
In broad terms, CoverUp contains components that perform three main functions: provide the LLM information on coverage, provide it with additional code context, and provide it with feedback on any errors.
In this research question, we explore how ablating each one of these components affects CoverUp.
Table 4 shows the specific elements disabled for each variant; disabling all three yields the same “CoverUp (ablated)” system explored already in RQ2.
Component
non-ablated
ablated [RQ2]
no coverage
no code context
no error fixing
Indication of what coverage is missing
Continued chat if generated tests do not improve coverage
Generated import statements
The get_info tool function
Continued chat in case of errors
Table 4. [RQ5] Elements in CoverUp ablations: The non-ablated CoverUp contains all of these elements, while CoverUp (ablated) from RQ2 contains none of them. To answer RQ5, we create three additional ablations.
Figure 17. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
Figure 18. [RQ5] Importance of coverage-based prompting: As the starting coverage grows with each successive run, the coverage-ablated prompt generates fewer and fewer functions that increase coverage.
We first evaluate these new ablations of CoverUp using suite CM; Figure 18 shows the resulting overall coverage for each ablation.
We observe that all ablations negatively affect coverage, indicating their importance.
Ablating error fixing has the most significant effect, losing between 14% and 37% on the various coverage metrics; ablating code context loses between 7% and 16%, and ablating coverage information up to 2%.
In interpreting these results, it is important to realize that the components are not entirely independent.
For example, the feedback from error fixing may allow the LLM to correct for insufficient code context, and through improved code context, errors resulting from incorrect assumptions may be avoided.
To better understand the effect of ablating coverage information from the prompt, we count the test functions generated by the model and assess how many of these increase coverage for each successive run.
As Figure 18 shows, as coverage increases, the coverage-ablated prompt results in fewer and fewer functions that increase coverage.
In contrast, CoverUp’s prompt remains effective.
This result shows that, by providing coverage information in the prompt, CoverUp focuses the LLM’s attention on uncovered code.
Even though the coverage-ablated prompt only fell short in coverage by a small amount, it did so by generating over 50% more functions than the non-ablated prompt (11,222 vs. 7,366).
If left in the test suite, these functions contribute to bloat, slowing the suite’s execution.
But even if rejected after a coverage check, they add unnecessary tokens in LLM responses and thus increase cost.
This analysis also suggests potential improvements for CoverUp.
Since even the non-ablated prompt results in some test functions that do not increase coverage, CoverUp should measure the coverage generated by each function individually, rather than for the entire LLM response.
Additionally, CoverUp could omit coverage information when prompting for tests for entirely uncovered functions, thereby reducing the cost of tagging every line in a code segment.
We leave this to future work.
Overall Coverage
Median Per-Module Coverage
Test Generator
@@ C. Tokens
Line + Branch
Line
Branch
Line + Branch
CoverUp (no coverage)
CoverUp (no code context)
CoverUp (no error fixing)
CoverUp (ablated) [RQ2]
Table 5. [RQ5] Effect of ablations on coverage: Removing any of these components from CoverUp lowers its performance.
First Run
Second Run
Third Run
End
Test Generator
InitialCov.
Cov. Incr.Functions
InitialCov.
Cov. Incr.Functions
InitialCov.
Cov. Incr.Functions
Cov.
TotalFunctions
CoverUp
7,366
CoverUp (no coverage)
11,222
Table 6. [RQ5] Importance of coverage-based prompting: As the starting coverage grows with each successive run, the coverage-ablated prompt generates fewer and fewer functions that increase coverage.
[RQ5] Summary: Coverage-based prompting, code context, and error fixing are all important components of CoverUp; ablating any of them results in reduced coverage.
5. Threats to Validity
Benchmark selection
We utilize the CM and PY benchmark suites to evaluate CoverUp’s performance on both challenging and less challenging code, as well as to facilitate including Codex-based results; we utilize the MT benchmark suite because MuTAP supports it, and it, too, facilitates including Codex-based results.
While our experience executing CoverUp to generate tests for other software has yielded similarly high coverage, selecting a different set of benchmarks could produce different results.
Execution environment
CodaMosa’s original evaluation environment failed to install a number of Python modules that are prerequisites for the applications used as benchmarks.
In an effort to replicate the original conditions as closely as possible, we ignored these failures as well.
It seems likely that both CodaMosa and CoverUp would be better able to generate tests if these modules were not missing.
LLM model dependency
CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo and GPT-4o models.
CoverUp’s approach is independent of the LLM and, as Section 4.3 shows, CoverUp significantly outperforms an ablated version of itself that just relies on the LLM’s capabilities.
Its ultimate performance naturally depends on the model’s ability to interpret its prompts and generate tests as requested.
6. Conclusion
This paper introduces CoverUp, a novel approach to guide LLM-based test generation through coverage analysis.
While modern LLMs achieve near-human performance on various tasks, simply prompting them is insufficient for generating high-coverage tests.
We demonstrate that integrating coverage information into prompts directs the LLM’s attention to uncovered code and highlight the benefits of iterative refinement through error and coverage feedback.
By combining these elements with code context, CoverUp produces test suites that achieve higher coverage than previous approaches.
Alshahwan et al. (2024)
Nadia Alshahwan, Jubin
Chheda, Anastasia Finogenova, Beliz
Gokkaya, Mark Harman, Inna Harper,
Alexandru Marginean, Shubho Sengupta,
and Eddy Wang. 2024.
Automated Unit Test Improvement using Large
Language Models at Meta. In Companion Proceedings
of the 32nd ACM International Conference on the Foundations of Software
Engineering, FSE 2024, Porto de Galinhas, Brazil, July 15-19, 2024,
Marcelo d’Amorim (Ed.).
ACM, 185–196.
https://doi.org/10.1145/3663529.3663839
Altmayer Pizzorno and Berger (2023)
Juan Altmayer Pizzorno and
Emery D. Berger. 2023.
SlipCover: Near Zero-Overhead Code Coverage for
Python. In Proceedings of the 32nd ACM SIGSOFT
International Symposium on Software Testing and Analysis (Seattle, WA, USA)
(ISSTA 2023). Association for
Computing Machinery, New York, NY, USA,
1195–1206.
https://doi.org/10.1145/3597926.3598128
arXiv:2305.02886
Bareiß et al. (2022)
Patrick Bareiß, Beatriz
Souza, Marcelo d’Amorim, and Michael
Pradel. 2022.
Code Generation Tools (Almost) for Free? A Study of
Few-Shot, Pre-Trained Language Models on Code.
arXiv:2206.01335 [cs.SE]
Boyapati et al. (2002)
Chandrasekhar Boyapati,
Sarfraz Khurshid, and Darko Marinov.
2002.
Korat: automated testing based on Java predicates.
In Proceedings of the 2002 ACM SIGSOFT
International Symposium on Software Testing and Analysis (Roma, Italy)
(ISSTA ’02). Association for
Computing Machinery, New York, NY, USA,
123–133.
https://doi.org/10.1145/566172.566191
Bubeck et al. (2023)
Sébastien Bubeck, Varun
Chandrasekaran, Ronen Eldan, Johannes
Gehrke, Eric Horvitz, Ece Kamar,
Peter Lee, Yin Tat Lee,
Yuanzhi Li, Scott Lundberg,
Harsha Nori, Hamid Palangi,
Marco Tulio Ribeiro, and Yi Zhang.
2023.
Sparks of Artificial General Intelligence: Early
experiments with GPT-4.
arXiv:2303.12712 [cs.CL]
Chen et al. (2021)
Mark Chen, Jerry Tworek,
Heewoo Jun, Qiming Yuan,
Henrique Ponde de Oliveira Pinto, Jared
Kaplan, Harri Edwards, Yuri Burda,
Nicholas Joseph, Greg Brockman,
Alex Ray, Raul Puri,
Gretchen Krueger, Michael Petrov,
Heidy Khlaaf, Girish Sastry,
Pamela Mishkin, Brooke Chan,
Scott Gray, Nick Ryder,
Mikhail Pavlov, Alethea Power,
Lukasz Kaiser, Mohammad Bavarian,
Clemens Winter, Philippe Tillet,
Felipe Petroski Such, Dave Cummings,
Matthias Plappert, Fotios Chantzis,
Elizabeth Barnes, Ariel Herbert-Voss,
William Hebgen Guss, Alex Nichol,
Alex Paino, Nikolas Tezak,
Jie Tang, Igor Babuschkin,
Suchir Balaji, Shantanu Jain,
William Saunders, Christopher Hesse,
Andrew N. Carr, Jan Leike,
Josh Achiam, Vedant Misra,
Evan Morikawa, Alec Radford,
Matthew Knight, Miles Brundage,
Mira Murati, Katie Mayer,
Peter Welinder, Bob McGrew,
Dario Amodei, Sam McCandlish,
Ilya Sutskever, and Wojciech Zaremba.
2021.
Evaluating Large Language Models Trained on Code.
arXiv:2107.03374 [cs.LG]
https://arxiv.org/abs/2107.03374
Chen et al. (2024)
Yinghao Chen, Zehao Hu,
Chen Zhi, Junxiao Han,
Shuiguang Deng, and Jianwei Yin.
2024.
ChatUniTest: A Framework for LLM-Based Test
Generation. In Companion Proceedings of the 32nd
ACM International Conference on the Foundations of Software Engineering,
FSE 2024, Porto de Galinhas, Brazil, July 15-19, 2024,
Marcelo d’Amorim (Ed.).
ACM, 572–576.
https://doi.org/10.1145/3663529.3663801
Daka and Fraser (2014)
Ermira Daka and Gordon
Fraser. 2014.
A Survey on Unit Testing Practices and Problems.
In 25th IEEE International Symposium on Software
Reliability Engineering, ISSRE 2014, Naples, Italy, November 3-6, 2014.
IEEE Computer Society, 201–211.
https://doi.org/10.1109/ISSRE.2014.11
Dakhel et al. (2024)
Arghavan Moradi Dakhel,
Amin Nikanjam, Vahid Majdinasab,
Foutse Khomh, and Michel C. Desmarais.
2024.
Effective test generation using pre-trained Large
Language Models and mutation testing.
Inf. Softw. Technol. 171
(2024), 107468.
https://doi.org/10.1016/J.INFSOF.2024.107468
Fioraldi et al. (2023)
Andrea Fioraldi,
Alessandro Mantovani, Dominik Christian
Maier, and Davide Balzarotti.
2023.
Dissecting American Fuzzy Lop: A FuzzBench
Evaluation.
ACM Trans. Softw. Eng. Methodol.
32, 2 (2023),
52:1–52:26.
https://doi.org/10.1145/3580596
Fraser and Arcuri (2011)
Gordon Fraser and Andrea
Arcuri. 2011.
EvoSuite: automatic test suite generation for
object-oriented software. In Proceedings of the
19th ACM SIGSOFT Symposium and the 13th European Conference on Foundations of
Software Engineering (Szeged, Hungary) (ESEC/FSE
’11). Association for Computing Machinery,
New York, NY, USA, 416–419.
https://doi.org/10.1145/2025113.2025179
Fraser and Arcuri (2014)
Gordon Fraser and Andrea
Arcuri. 2014.